# EmotiCrafter: Text-to-Emotional-Image Generation based on Valence-Arousal Model

### v1

@@ Abstract
Recent research shows that emotions can enhance users’ cognition and influence information communication. While research on visual emotion analysis is extensive, limited work has been done on helping users generate emotionally rich image content. Existing work on emotional image generation relies on discrete emotion categories, making it challenging to capture complex and subtle emotional nuances accurately. Additionally, these methods struggle to control the specific content of generated images based on text prompts. In this work, we introduce the new task of continuous emotional image content generation (C-EICG) and present EmotiCrafter, an emotional image generation model that generates images based on text prompts and Valence-Arousal values. Specifically, we propose a novel emotion-embedding mapping network that embeds Valence-Arousal values into textual features, enabling the capture of specific emotions in alignment with intended input prompts. Additionally, we introduce a loss function to enhance emotion expression. The experimental results show that our method effectively generates images representing specific emotions with the desired content and outperforms existing techniques.
@@
Yi He and Shengqi Dang contributed equally to this work.Nan Cao is the corresponding author.
@@ 1 Introduction
Emotions are fundamental to human experiences and play a critical role in shaping how people perceive and interact with the world. Research has shown that emotions affect memory [24, 20, 30, 5, 47] and comprehension [21, 42, 40], which are crucial for effective communication. As a result, content creators increasingly recognize the importance of incorporating emotions to enhance audience engagement and comprehension.
@@
While research on visual emotion analysis is extensive [52, 18, 61], there is limited work on generating emotionally rich image content. Some early studies explored emotion generation techniques within specific domains such as facial expressions [2, 48] or landscapes [29]. EmoGen [53] generates images based on a given emotion tag (e.g., happy or sad), which pioneered the domain-free emotional image generation (EICG) task. However, it has two critical limitations: (1) the image is generated from an emotion tag instead of a text prompt, making the generated content difficult to control; (2) While the discrete emotion tags used in EmoGen are easy to understand, psychologists have not achieved a consensus on the emotion categories [63]. The limited scope of discrete emotion tags falls short of capturing nuanced emotions. In contrast, the continuous emotion models represent emotions through a continuous Cartesian coordinate system, allowing for a more precise capture of emotional complexity and subtlety [2]. Therefore, developing a general text-to-image model that reflects continuous emotional variations is desired.
In this paper, we introduce the task of continuous emotional image content generation (C-EICG) and present EmotiCrafter, an emotional image generation model that generates images based on text prompts and V-A values (i.e., two real numbers $(v,a)\in[-3,+3]$) defined in the Valence-Arousal (V-A) emotion model [37], the most frequently used continuous emotional model in psychology. An emotion-embedding network is designed to embed V-A values into the text prompts of a stable diffusion XL model (SDXL) [32]. This network enables a nonlinear embedding of V-A values into the features of the input text prompt to capture complex emotional context for image generation. To train the model, a dataset with 39,843 annotated images is prepared. Given an image, a training sample consists of a pair of V-A values about the image labeled by human annotators and two prompts, i.e., neutral prompt and emotional prompt generated by GPT-4 based on the image. The neutral [emotion] prompt describes the essential content of the image, whereas the emotional prompt describes the emotional information, such as the color, light, and texture, about the image. Finally, to enhance the generation results, we introduce a loss function that normalizes the V-A distribution and scales the residuals between the features of the neutral and emotional prompts provided in each training sample. In summary, our contributions are as follows:
  • We introduce a new task, i.e., continuous emotional image content generation (C-EICG), and propose a novel emotion-embedding network based on the transformer architecture. The network embeds the continuous Valence-Arousal values into the textual features of an input prompt for emotional image generation.
  • We develop a novel loss function that helps to enhance the emotional resonance of generated images. The loss function modifies the target features to learn more obvious emotional changes and takes the V-A distribution into account to counter the unevenness of the data.
@@
  • We have constructed an emotional image dataset for training the proposed emotion-embedding network. Each training sample consists of a pair of unified V-A values, a neutral prompt, and an emotional prompt that are associated with the same image in the dataset.
@@ 2.1 Visual Emotion Analysis
Much of the continuous emotion analysis remains centered on facial expression analysis [39, 27, 57, 13, 15]. While effective, facial analysis can overlook crucial contextual information that influences emotional interpretation. This drives studies to focus on emotions in objects or individuals within their environment, rather than just faces[26, 26, 16, 25, 18, 18]. For example, Kosti et al. [16] combined person-specific characteristics with the context of the scene to predict continuous emotional dimensions such as arousal, valence, and dominance. Kragel et al. [18] proposed EmoNet to extract visual features such as facial expressions, body posture, and scene elements from images to predict valence and arousal values. Recently, Mertens et al.[26] tested multiple backbones (such as ResNet, CLIP, DINO, etc.) for prediction of valence and arousal, demonstrating strong performance across various architectures.
Previous studies have achieved high accuracy in predicting continuous emotions in images, demonstrating a correlation between continuous emotions and visual elements. Building on this insight, our work moves beyond prediction to explore how continuous emotions can be actively embedded within generated content. Specifically, we introduce the task of continuous emotional image content generation (C-EICG) and propose a novel emotion-embedding network.
@@ 2.2 Image Emotion Transfer
Image Emotion Transfer (IET) focuses on editing the content of the images to evoke different emotions [22, 64, 65, 31, 45]. For instance, Peng et al. [31] achieved emotion transfer by adjusting color tones and texture-related features. Zhu et al. [65] introduced a method that separates high-level emotion-relevant features (e.g., object shapes and scene layout) from low-level emotion-relevant features (e.g., brightness). By applying GANs, they transferred emotions between images while preserving their original structure. Building on these methods, Weng et al. [45] proposed the Affective Image Filter, which uses a multi-modal transformer to process both text and image inputs. With the emergence of text-to-image models, IET has expanded into new applications based on instructive commands. For example, EmoEdit [54] used GPT-4V to build emotion factor trees that map abstract emotions to specific visual elements and employed the InstructPix2Pix model to apply emotion-driven content and color adjustments to images.
Current methods primarily extract specific features, often focusing on certain visual elements or emotional cues, which can limit the depth of emotional expression. In contrast, our method broadly learns a variety of emotion-influencing features through the emotion-embedding network, mapping continuous emotional variations at the textual level. Additionally, our model can accept natural language prompts as input, allowing for a more intuitive and flexible way to generate emotionally expressive images.
@@ 2.3 Conditional Image Generation
Conditional image generation aims to produce images that align with specific input conditions, such as text, depth maps, or sketches. Early explorations in this field utilized Variational Autoencoders (VAEs) [14] and Generative Adversarial Networks (GANs) [9]. Currently, textual conditional image generation models based on denoising diffusion [12], such as Stable Diffusion [35, 32] and DALL-E [34, 3], can generate exceptionally high-quality images. Besides, several studies have explored the use of other pieces of information as conditions for generation. The IP-adapter [56] integrates features from reference images through cross-attention, generating images consistent with attributes or identity. Techniques like Textual Inversion [8] and DreamBooth [36] can generate images in specific styles or domains by learning a concept. Models like ControlNet [58] and T2I-Adaptor [28] facilitate flexible conditional control using sketches, depth maps, and others.
However, there is limited research on the use of emotion as a condition. EmoGen [53] pioneered the generation of emotional image content (EIGC) by mapping emotional features to semantic features to create emotional images. Despite this, EmoGen struggles to understand natural language, limiting its capacity to effectively control specific content. Its reliance on discrete emotions also limits its practical applicability.
Our method bridges this gap by embedding continuous emotions into textual features, enabling image generation models to use both emotion and input prompts as conditions for creating emotional images.
@@
Figure 3: Structure of Emotion Injection Block. It accepts hidden state $h_{i-1}$ as input and produces $h_{i}$ as output. The V-feature $e_{v}$ and A-feature $e_{a}$ represent the emotional features, which are injected through the cross-attention module.
@@ 3 Method
In this section, we introduce the technique details of EmotiCrafter. We begin with an overview of the proposed method, followed by an explanation of the network structure and key modules. Subsequently, we outline the training strategy. Finally, we describe the process of collecting and annotation of our dataset for network training.
@@ 3.1 Overview
As illustrated in Figure 2, our method generates emotional images $I_{emo}$ from two inputs: text prompts describing the desired image content and a pair of V-A values $(v,a)$ specifying the emotions. The process begins by encoding the input prompts into the prompt features $f_{n}$ using a prompt encoder $\mathcal{E}$. This features then pass through an emotion-embedding network $\mathcal{M}$ (Figure 2 (b.1)) to produce the emotional prompt features $\hat{f}_{e}$, embedded with emotion values $(v,a)$:
$\hat{f}_{e}$ $=\mathcal{M}(f_{n}|(v,a)).$ (1)
Finally, the emotional prompt features are input via the cross-attention mechanism in the SDXL $\mathcal{G}$ to generate the emotional images $I_{emo}=\mathcal{G}(\hat{f}_{e})$. To enhance the emotional expressiveness of generated images, we develop a loss function (Figure 2 (b.2)) that optimizes the network training by considering the V-A distribution of the training data and enhancing the residuals between the features of the emotional prompts and neutral prompts, compelling that generated images convey the desired emotions and content. To facilitate our emotion-embedding network, we have compiled an image dataset (Figure 2 (a)). This dataset pairs emotional images with both neutral prompts and emotional prompts, along with corresponding V-A values.
@@ 3.2 Emotion-Embedding Network
The emotion-embedding network $\mathcal{M}$ takes a pair of V-A values and a neutral prompt feature as input and outputs the corresponding emotional prompt feature (Figure 2 (b.1)). In particular, it first employs a V-A encoder to convert the V-A values into feature vectors. After that, an emotion injection transformer network (structure modified from GPT-2 [33] in our implementation) is used to infuse these features into the neutral prompt feature. Here, the transformer helps with capturing and preserving the textual context of the neutral prompt while injecting emotional features.
V-A Encoder. The V-A Encoder transforms a pair of V-A values into emotional features using two separate multilayer perceptrons (MLPs). The output V-features $e_{v}$ and the A-features $e_{a}$ are then transmitted into the emotion infusion process of the emotion injection transformer network.
Emotion Injection Transformer (EIT). The EIT is designed to effectively embed V/A-features into textual features through a modified GPT-2 architecture. The computing process of EIT consists of three steps: input projection, emotion injection, and output projection. First, we project the input neutral prompt feature $f_{n}$ into the transformer’s feature space:
@@
where $h_{0}$ represents the initial hidden state, $\text{P}_{\text{in}}(\cdot)$ is a linear projection layer, and PE denotes positional embedding [41] The core of EIT consists of 12 sequential Emotion Injection Blocks (EIBs) corresponding to 12 transformer blocks for emotion injection. Each block outputs a hidden state:
@@
where $h_{i}$ is the output of the $i$-th EIB()($\cdot$)( ⋅ ). As shown in Figure 3, each EIB enhances the traditional transformer block through a cross-attention mechanism:
$h^{\prime}_{i}=\text{self-attn}(\text{LN}(h_{i-1}))+h_{i-1},$ (4)
$h^{(v)}_{i}=\text{cross-attn}(\text{LN}(h^{\prime}_{i}),e_{v})+h^{\prime}_{i},$ (5)
$h^{(v,a)}_{i}=\text{cross-attn}(\text{LN}(h^{(v)}_{i}),e_{a})+h^{(v)}_{i},$ (6)
$h_{i}=\text{fnn}(\text{LN}(h^{(v,a)}_{i}))+h^{(v,a)}_{i}),$ (7)
@@
$\hat{f}_{r}=\text{P}_{\text{out}}(\text{LN}(h_{12})),$ (8)
@@ 3.4 Dataset and Training
To train the emotion embedding network, an image caption dataset built on a set of annotated images is prepared. As shown in Figure 2, given an image, a sample in our dataset is a 3-tuple that consists of (1) a pair of V-A values about the image labeled by human annotators and two prompts, i.e., (2) neutral prompt and (3) emotional prompt. Both prompts are automatically generated using GPT-4, taking the image as the input. In particular, the neutral [emotion] prompt describes the essential content of the image, whereas the emotional prompt describes the emotional information, such as the color, light, and texture, about the image.
In particular, we collected OASIS [19], EMOTIC [17], and FindingEmo [26] to gather enough images with human-annotated V-A values. These images offer a diverse range of scenes beyond just facial images. After cleaning and merging, we compiled a final dataset of 39,843 images with V-A annotations. We applied linear scaling to standardize the varying V-A scales across these datasets, adjusting the values to a uniform range of $[-3,3]$ [26, 50]. Next, we applied prompt engineering to annotate the processed datasets with both neutral and emotional prompts, leveraging GPT-4’s vision capabilities to recognize essential content and express emotional nuances [7]. Please refer to the supplementary materials for further details.
The proposed emotion embedding network is trained based on two NVIDIA A800 GPUs using the above dataset. We utilize the AdamW [23] optimizer with a weight decay of 1e-5 and a learning rate of 1e-3. The training lasted for approximately 7 to 8 hours over a total of 200 epochs and 768 batch-size.
4 Experiments
In this section, we present both qualitative and quantitative comparisons, along with a user study and an ablation study, to evaluate the effectiveness of the proposed method.

### v2

@@ Abstract
Recent research shows that emotions can enhance users’ cognition and influence information communication. While research on visual emotion analysis is extensive, limited work has been done on helping users generate emotionally rich image content. Existing work on emotional image generation relies on discrete emotion categories, making it challenging to capture complex and subtle emotional nuances accurately. Additionally, these methods struggle to control the specific content of generated images based on text prompts. In this paper, we introduce the task of continuous emotional image content generation (C-EICG) and present EmotiCrafter, a general emotional image generation model that generates images based on free text prompts and Valence-Arousal (V-A) values. It leverages a novel emotion-embedding mapping network to fuse V-A values into textual features, enabling the capture of emotions in alignment with intended input prompts. A novel loss function is also proposed to enhance emotion expression. The experimental results show that our method effectively generates images representing specific emotions with the desired content and outperforms existing techniques.
@@
00footnotetext: Shengqi Dang and Yi He contributed equally to this work.00footnotetext: Nan Cao is the corresponding author.00footnotetext: Shengqi Dang, Yi He, Long Ling, Ziqing Qian, and Nan Cao are with the Intelligent Big Data Visualization Lab, Tongji University. Shengqi Dang and Nan Cao are also with the Shanghai Innovation Institution. Email: {dangsq123, heyi_11}@tongji.edu.cn, lucyling0224@gmail.com, 2411920@tongji.edu.cn, nan.cao@gmail.com.00footnotetext: Nanxuan Zhao is with Adobe Research. Email: nanxuanzhao@g- mail.com.00footnotetext: Code: https://github.com/idvxlab/EmotiCrafter
@@ 1 Introduction
Emotions are fundamental to human experiences and play a critical role in shaping how people perceive and interact with the world. Research has shown that emotions affect memory [19, 15, 25, 45] and comprehension [16, 40, 38], which are crucial for effective communication. As a result, content creators increasingly recognize the importance of incorporating emotions to enhance audience engagement.
@@
While research on visual emotion analysis is extensive [49, 13, 58], there is limited work on generating emotionally rich image content. Some early studies explored emotional content generation techniques within specific domains such as facial expressions [2, 46] or landscapes [24]. EmoGen [50] generates images based on a given emotion tag (e.g., happy or sad), which pioneered the domain-free emotional image content generation (EICG) task. However, it has two critical limitations: (1) the image is generated from an emotion tag instead of a text prompt, making the generated content difficult to control; (2) while the discrete emotion tags used in EmoGen are easy to understand, psychologists have not achieved a consensus on the emotion categories [60]. The limited scope of discrete emotion tags falls short of capturing nuanced emotions.
To address the above issues, we propose the continuous emotional image content generation (C-EICG) task and present EmotiCrafter, the first C-EIGC model that generates emotional images using free-text prompts and continuous Valence-Arousal (V-A) values defined in the V-A model (a well-known psychological continuous emotion model)  [33]. V-A model represents emotions in a two-dimensional Cartesian space (Figure 2), where Valence quantifies pleasantness (negative to positive) and Arousal measures intensity (calm to excited). We utilize a [-3, 3] range for Valence and Arousal [10, 21]. The continuous V-A space enables smooth transitions and nuanced emotional shifts beyond the capability of discrete labels (e.g., shifts from “bored” to “tired” of a character’s state in a video). Specific emotional values in this space have been investigated in prior work [34], and such fine-grained modeling is particularly beneficial for human–computer interaction [2]. Leveraging this model, EmotiCrafter captures subtle affective variations via precise (V, A) positioning. The contributions of this paper are as follows:
Figure 2: Valence-Arousal model.
  • We propose a novel task, continuous emotional image content generation (C-EICG), and develop the first dedicated model for this purpose. Our model introduces an emotion-embedding network that integrates continuous Valence-Arousal (V-A) values with text prompts. These fused features are then injected into Stable Diffusion XL [28] using cross-attention mechanisms, enabling precise control over both content and emotional expressions in the generated images.
  • We propose a novel loss function that enhances the emotional resonance of generated images. By amplifying the difference between neutral and emotional text features, our approach enables the model to capture more distinct emotional variations. Additionally, the loss function incorporates the V-A distribution to address data imbalance, further refining the model’s ability to generate images with rich, accurate emotional expressions.
@@
  • We constructed an emotional prompts dataset to train the emotion-embedding network, where each sample consists of a neutral prompt and an emotional prompt that share the same core meaning but express a specific emotion corresponding to a given pair of V-A values.
@@ 2.1 Visual Emotion Analysis
Much of the continuous emotion analysis remains centered on facial expression analysis [37, 22, 54, 9, 11]. While effective, facial analysis can overlook crucial contextual information that influences emotional interpretation. This drives studies to focus on emotions in objects or individuals within their environment, rather than just faces[21, 12, 20, 13].
For example, Kosti et al. [12] combined person-specific characteristics with the context of the scene to predict continuous emotional dimensions such as arousal, valence, and dominance. Kragel et al. [13] proposed EmoNet to extract visual features such as facial expressions, body posture, and scene elements from images to predict valence and arousal values. Recently, Mertens et al.[21] tested multiple backbones (such as ResNet, CLIP, DINO, etc.) for prediction of valence and arousal, demonstrating strong performance across various architectures.
Previous studies have achieved high accuracy in predicting continuous emotions in images, demonstrating a correlation between continuous emotions and visual elements. Building on this insight, our work moves beyond prediction to explore how continuous emotions can be actively embedded within generated content.
@@ 2.2 Image Emotion Transfer
Image Emotion Transfer (IET) focuses on editing the content of the images to evoke different emotions [17, 61, 27, 43]. For instance, Peng et al. [27] achieved emotion transfer by adjusting color tones and texture-related features. Zhu et al. [61] introduced a method that separates high-level emotion-relevant features (e.g., object shapes and scene layout) from low-level emotion-relevant features (e.g., brightness). By applying GANs, they transferred emotions between images while preserving their original structure. Building on these methods, Weng et al. [43] proposed the Affective Image Filter, which uses a multi-modal transformer to process both text and image inputs. With the emergence of text-to-image models, IET has expanded into new applications based on instructive commands. For example, EmoEdit [51] used GPT-4V to build emotion factor trees that map abstract emotions to specific visual elements and employed the InstructPix2Pix model to apply emotion-driven content and color adjustments to images.
Current methods primarily extract specific features, focusing on certain visual elements or emotional cues, which can limit the depth of emotional expression. In contrast, our method broadly learns a variety of emotion-influencing features, and it could accept natural language prompts as input.
@@ 2.3 Conditional Image Generation
Conditional image generation aims to create images that align with specific input conditions, such as text [31, 28, 30, 3], reference images [53], subjects [6, 32], and depth maps [55, 23]. To enhance quality, researchers have developed specialized approaches, such as Diffusion Transformers (DiT)[26], which use transformer-based diffusion for denoising, and Visual Autoregressive Modeling (VAR)[36], which encodes images into discrete tokens for autoregressive prediction across scales. However, these state-of-the-art methods rely on discrete labels, limiting flexibility and control. In contrast, our approach offers greater generalization by leveraging Stable Diffusion XL [28] to generate emotional images from free-text prompts.
Despite these progresses, incorporating emotion as a condition for image generation remains underexplored. EmoGen [50] pioneered the generation of emotional image content (EIGC) by mapping emotional features to semantic features to generate emotional images. However, EmoGen struggles to understand natural language, limiting its capacity to effectively control specific content. Its reliance on discrete emotions also limits its practical applicability.
Our method bridges this gap by embedding continuous emotion into textual features, enabling image generation models to use continuous emotion for emotion control. Unlike label-based methods, our method supports free-text prompts for flexible content control.
@@
Figure 4: Structure of Emotion Injection Block. It accepts hidden state $h_{i-1}$ as input and produces $h_{i}$ as output. The V-feature $e_{v}$ and A-feature $e_{a}$ represent the emotion features, which are injected through the cross-attention module.
@@ 3 Method
In this section, we introduce the technique details of the proposed EmotiCrafter.
@@ 3.1 Overview
Our method generates emotional images $I_{emo}$ from two inputs (Figure 3): a free-text prompt describing the desired content, and a pair of V-A values $(v,a)$ specifying the emotion. First the prompt encoder $\mathcal{E}$ converts the text prompt into feature $f_{n}$, which is then processed by an emotion-embedding network $\mathcal{M}$ to produce emotional prompt feature $\hat{f}_{e}$ that integrate the V-A values:
$\hat{f}_{e}$ $=\mathcal{M}(f_{n}|(v,a))$ (1)
Next, this feature is injected into Stable Diffusion XL $\mathcal{G}$ via its cross-attention mechanism to generate the emotional images: $I_{emo}=\mathcal{G}(\hat{f}_{e})$. To enhance emotional expressiveness, we introduce a loss function (Fig.3(b.2)) that leverages the V-A distribution and emphasizes the differences between neutral and emotional prompt features. This ensures that the generated images accurately convey both the intended emotions and content. Additionally, we construct a dataset (Fig.3(a)) that pairs neutral and emotional prompts with the corresponding V-A values.
@@ 3.2 Emotion-Embedding Network
The emotion-embedding network $\mathcal{M}$ generates the emotional prompt feature by integrating a pair of V-A values with a neutral prompt feature (Fig.3(b.1)). First, a V-A encoder converts the V-A values into feature vectors. Then, an emotion injection transformer—modified from GPT-2[29]—fuses these vectors with the neutral prompt feature, preserving the original textual context while infusing emotional content.
V-A Encoder. The V-A Encoder converts a pair of V-A values into feature vectors using two separate multilayer perceptrons (MLPs). One MLP processes the Valence value to produce V-feature $e_{v}$, and the other processes the Arousal value to generate A-feature $e_{a}$. These features are then fed into the emotion injection transformer network for further emotion infusion.
Emotion Injection Transformer (EIT). The Emotion Injection Transformer (EIT) leverages a modified GPT-2 architecture to seamlessly integrate V/A-features into textual features. Its process consists of three stages: input projection, emotion injection, and output projection.
First, we project the input neutral prompt feature $f_{n}$ into the transformer’s feature space:
@@
where $h_{0}$ represents the initial hidden state; $\text{P}_{\text{in}}(\cdot)$ is a linear projection layer, and PE denotes positional embedding [39].
In the next, we inject emotion into $f_{n}$ via 12 sequential Emotion Injection Blocks (EIBs) corresponding to 12 transformer blocks. Each block outputs a hidden state:
@@
where $h_{i}$ is the output of the $i$-th EIB()($\cdot$)( ⋅ ). As shown in Figure 4, each EIB enhances the transformer block through a cross-attention mechanism:
$h^{\prime}_{i}=\text{self-attn}(\text{LN}(h_{i-1}))+h_{i-1}$ (4)
$h^{(v)}_{i}=\text{cross-attn}(\text{LN}(h^{\prime}_{i}),e_{v})+h^{\prime}_{i}$ (5)
$h^{(v,a)}_{i}=\text{cross-attn}(\text{LN}(h^{(v)}_{i}),e_{a})+h^{(v)}_{i}$ (6)
$h_{i}=\text{fnn}(\text{LN}(h^{(v,a)}_{i}))+h^{(v,a)}_{i}$ (7)
@@
Finally, the output of the last (i.e., the 12th) injection block $h_{12}$ is projected back to SDXL’s prompt feature space via $\text{P}_{\text{out}}$ (a linear layer and a LayerNorm layer) to predict the residual between emotional and neutral prompt features, which represents a semantic shift between emotional and neutral prompts:
$\hat{f}_{r}=\text{P}_{\text{out}}(\text{LN}(h_{12}))$ (8)
@@ 3.3 Dataset and Training
To train the emotion-embedding network, we constructed a dataset of paired neutral and emotional prompts with corresponding Valence-Arousal (V-A) values (Figure 3(a)). These pairs were automatically generated using GPT-4 based on 39,843 images with human-annotated V-A values from publicly available datasets, including OASIS [14], EMOTIC [12], and FindingEmo [21]. Specifically, GPT-4 was used to generate neutral prompts with objective image descriptions and emotional prompts that emphasize affective attributes such as color, lighting, and texture, which influence emotional perception. To ensure data reliability, all LLM-generated prompts underwent crowd-sourced verification, with disagreements resolved through a voting mechanism among annotators.
The proposed emotion embedding network is trained on two NVIDIA A800 GPUs using the aforementioned dataset. We employ the AdamW [18] optimizer with a weight decay of 1e-5 and a learning rate of 1e-3. The training process spans 200 epochs with a batch size of 768, completing in approximately 7 to 8 hours.
Refer to caption (a)Refer to caption (b)Refer to caption (c)Refer to caption (d)
Figure 5: Results under multiple inputs. (a) Overriding semantic content (‘a child in the amusement park’) with sad V-A (-2,-2); (b) Discrete emotion mapping in V-A space as emotion input; (c) Empty-prompt generation with pure emotion condition; (d) Fine-grained control of V-A variations with a granularity of 0.2.
4 Evaluation